Activity of a radioactive substance decreases from 8000BQ ato 1000Bq in 9 days.What is half life and average life of radoactive substance

Grade:12

1 Answers

Chandra Chatterjee
9 years ago
 
Let A be the new activity and A’ be the original activity
Now there is a formula (A/A’) = (1/2)^(t/T). Where t is the given time and T is the half life.  (I am not showing the derivation. It will become long. U can refer to any good book)
Now it's given A=1000 A’=8000 t=9.  Therefore (1000/8000)=(1/2)^9/T  
Or, (1/2)^3=(1/2)^9/T.     So 3= 9/T So T=3.        Then comes average life.      Here also there is a simple relation between average life T.. and half life T  that is T..= 1.44T.   So avg life is 4.32days.